OpenStudy (superhelp101):

Which of the following is a possible set of quantum numbers for an electron n, l, m subscript l, m subscript ? (1, 0, 0, +1) (2, 2, 1, +½) (4, -3, 0, -½) (3, 2, 1, -½)

OpenStudy (superhelp101):

@Abhisar is this like the first problem we did?

OpenStudy (abhisar):

What is n, l, m subscript l, m subscript ?

OpenStudy (abhisar):

is it n, l, \(\huge\text{m}\)\(\huge_l\), \(\huge\text{m}_s\) ?

OpenStudy (abhisar):

if yes then yes it's like the first question

OpenStudy (abhisar):

Let's see if (2, 2, 1, +½) is possible or not

OpenStudy (abhisar):

For n=2, l is 0 to (n-1) l = 0,1 (l can never be equal to n)

OpenStudy (abhisar):

So the option is incorrect !

OpenStudy (abhisar):

Check the first option

OpenStudy (superhelp101):

ok

OpenStudy (superhelp101):

yes it works 0 to (1-1) 0=0

OpenStudy (abhisar):

Hint: \(\huge\text{m}_s\) has only two values +1/2 and -1/2

OpenStudy (abhisar):

yes it works for l but does it works for \(\huge\text{m}_s\) too ? check it !

OpenStudy (abhisar):

u will have to check each quantum number in a given option

OpenStudy (abhisar):

Lets see one more example (4, -3, 0, -½)

OpenStudy (abhisar):

do with me

OpenStudy (superhelp101):

ok

OpenStudy (abhisar):

Lets first check l

OpenStudy (abhisar):

n=4, so l = ?

OpenStudy (superhelp101):

0 to (4-1) 0 to 3 i.e 1,2,3

OpenStudy (abhisar):

yes i.e 0,1,2,3 ryt ?

OpenStudy (superhelp101):

oh yes

OpenStudy (abhisar):

now check if the option says the same ?

OpenStudy (superhelp101):

no

OpenStudy (abhisar):

very good !

OpenStudy (abhisar):

it says -3, so the option is wrong

OpenStudy (abhisar):

ok ?

OpenStudy (superhelp101):

ok

OpenStudy (abhisar):

Now Lets check (1, 0, 0, +1)

OpenStudy (abhisar):

n=1, so l= ?

OpenStudy (superhelp101):

they all work

OpenStudy (abhisar):

wait, first tell me l=?

OpenStudy (superhelp101):

0 to (n-1) 0 to (1-1) 0=0 0 to (1-1) 0=0 and what is the difference between 1 and +1

OpenStudy (abhisar):

no difference between +1 and 1, both are same but -1 & 1 or -1 & +1 are different

OpenStudy (abhisar):

ohkay so l will have only one value this time and that is 0

OpenStudy (superhelp101):

yes

OpenStudy (abhisar):

even the option says 0, so it is correct upto l. Now lets move to \(\huge\text{m}_l\) in the same option

OpenStudy (superhelp101):

wait so are you saying that the first option is incorrect?

OpenStudy (abhisar):

no i am just saying that it is correct upto l. To say that it is completely correct we will have to check the rest two quantum numbers too

OpenStudy (abhisar):

getting it ?

OpenStudy (superhelp101):

oh yea ok

OpenStudy (superhelp101):

so are we doing the second option now?

OpenStudy (abhisar):

no u have to check the rest quantum numbers in the first option. Check if \(\huge m_s\) is correct or not

OpenStudy (superhelp101):

oh ok

OpenStudy (abhisar):

\(\huge m_s\) = -l to +l

OpenStudy (abhisar):

we got l=0 ryt ?

OpenStudy (superhelp101):

yes

OpenStudy (superhelp101):

so that can't be right

OpenStudy (abhisar):

so \(\huge m_s\) will be -0,0 &+0 or simply say only 0

OpenStudy (abhisar):

getting it ?

OpenStudy (superhelp101):

ooo yes

OpenStudy (abhisar):

Option A also says that \(\huge m_s\) = 0, so we know now that 3 out of four quantum numbers are correct, lets check the final one

OpenStudy (superhelp101):

alright

OpenStudy (abhisar):

oh man ! Sorry \(\huge m_s\) will be \(\huge m_l\) in my last 4 replies. i typed it wrong. Please correct it

OpenStudy (superhelp101):

yea I was thinking that

OpenStudy (abhisar):

now lets check the fourth one which is \(\huge m_s\)

OpenStudy (abhisar):

\(\huge m_s\) has only 2 values (always), either +1/2 or -1/2. check what is given in option A

OpenStudy (superhelp101):

0 to (1/2)-1) 0 to 1-(1/2) 0 to 1/2

OpenStudy (abhisar):

what r u doin ? Just check the 4th quantum number in option A

OpenStudy (abhisar):

Simply see if it is one among +1/2 or -1/2 or not

OpenStudy (superhelp101):

Huh? like this. 0 to (1-1) 0=0

OpenStudy (abhisar):

u r not getting the pont....each quantum number has its own way of calculation. Formula for calculating every quantum number is different. There are four different ways to calculate the four different quantum numbers

OpenStudy (superhelp101):

ooo

OpenStudy (abhisar):

(1, 0, 0, +1), this means 4 different quantum numbers

OpenStudy (abhisar):



OpenStudy (superhelp101):

It is among 1/2 but not -1/2

OpenStudy (abhisar):

u have quantum numbers given in ur textbook ?

OpenStudy (superhelp101):

huh?

OpenStudy (abhisar):

read it ! I will see u tmrw ! it's 2:20 AM here

OpenStudy (superhelp101):

oh ok thanks thanks! what country are u in?

OpenStudy (superhelp101):

@Abhisar

OpenStudy (abhisar):

http://openstudy.com/users/abhisar

OpenStudy (superhelp101):

India!